OpenStudy (cryleeb):

WILL FAN AND MEDAL A student is trying to solve the system of two equations given below: Equation P: y + z = 6 Equation Q: 5y + 9z = 1 Which of the following is a possible step used in eliminating the y-term? (1 point)

OpenStudy (anonymous):

Options?

OpenStudy (cryleeb):

a. (y + z = 6) ⋅9 b. (y + z = 6) ⋅–5 c. (5y + 9z = 1) ⋅9 d. (5y + 9z = 1) ⋅5

OpenStudy (cryleeb):

@Zelda, do you think you can help me with this? (:

OpenStudy (anonymous):

Well, if you wanted to cancel out the y terms, they'd have to add up to zero, right? So which option would make the top equation have -5y?

OpenStudy (cryleeb):

B?